QUESTION 2 For the below loaded truss, P1-20 kN, P2 30 kN, P3-40 kN. 3 m B 12 m--2 m--2 m--2 m--2 m-1-2 m- P Choose the best answer for the force in member KC in kN if the vertical reaction Gy 33.33 kN. State d the member is in tension (T) or compression (C) 12.67 (C) a.67 (C) 6.6 (D 12.67 (
For the below loaded truss, P1-20 kN, P2 30 KN, P3-40 kN. H 3 m D F -2 m--2 m--2 m-1-2 m-1-2 m- P P: Choose the best answer for the force in member CD in kN. if the vertical reaction Gys33 33 kN State if the member is in tension (T) or compression (C). 31.11 (T) None of the answers 31.11 (C) 62 22 (T 62 22 (C)
Answer #1

For the below loaded truss, P1-20 kN, P2 30 kN, P3-40 kN. K H 3 m C E F E -2 m--2 m--2 m--2 m-1-2 m-1-2 m- n--2 m-1-2 m-1-2 m

consider the portion of the truss to the right of the section a-a, shown above

Darctan(3/2)= 56.31

EM 033.33 6KJ3 0 KJ=66.66kN (c)

EF,0 33.33 40 KD sin 56.31 0 KD = 7/sin 56.31 8.017kN (T)

EF, 0 -KJ-KD cos 56.31 CD = 0

66.66 8.017 cos 56.31°-CD = 0 CD -62.22kN(C) answer

QUESTION 2 KL , P1-20 kN, P2-30 KN, P3-40 kN. For the below loade K H K KC 3 m E CD E B C F +2 m--2 m--2 m- -2 m--2 m--2 m--2

consider the portion of the truss to the right of the section b-b

EF 0 KC 4033.33 0 KC = -6.67kN(C) answer